OpenStudy (anonymous):

What is the inverse function of f(x) = 4x + 3?

OpenStudy (loser66):

let y = 4x +3 switch y by x x = 4y +3 then solve for "new" y 4y = x-3 ---> y = (x-3)/4 that "new" y is inverse of the original one. using notation